OpenStudy (anonymous):

Ok, I have another one... How do I solve this - Use prime factors to show that 3136 is a perfect square... how do I break it down and teach this

OpenStudy (anonymous):

3136 = 2*2*2*2*2*2*7*7 = (2*2*2*7)*(2*2*2*7) = 56 * 56